# Multi-scale analysis of flow over heterogeneous urban environments

### v1

@@ 4 Results
4.1 Classical analysis
@@
Figure 3: Time-averaged flow statistics. (a) Wind speed $\sqrt{\overline{u}^{2}+\overline{v}^{2}}$ at mean building height $z=30$ m overlaid with $(\overline{u},\overline{v})$ velocity vectors; (b) wind speed $\sqrt{\overline{u}^{2}+\overline{w}^{2}}$ at the central vertical plane $y=384$ m overlaid with ($\overline{u}$,$\overline{w}$) velocity vectors; (c) Turbulent kinetic energy $k$ at mean building height $z=30$ m. The white boxes represent the buildings.
@@
The time-averaged wind speed at the mean building height $z=30$ m (figure 3a) shows large variations in wind speed and channelling in regions where buildings are absent. A strong reduction in wind speed is shown near some of the tall building clusters (Mishra et al. 2023, 2024), with wakes clearly visible downstream of the tall buildings. The elevation view (figure 3b) shows the general increase of wind speed with height and also demonstrates the homogenisation of the wind in the $x$ direction with increasing height, consistent with the extent of the roughness sublayer (the height of this layer is $h_{\max}$ as will be shown in figure 4). The velocity vectors show a clearly visible wake behind the building (in particular behind the first building in figure 3b). Accelerations are visible near the upwind top of the (e.g., for the second cube in figure 3b), which is consistent with Coceal et al. (2007). Also visible are skimming flows, e.g., near the last cube in the figure; and the flow circulation, e.g., between the last third and last second cube in figure 3(b).
@@
Figure 3(c) shows the plane view at $z=30$ m of time-averaged turbulent kinetic energy (TKE, $k$). Note that the TKE field is correlated with the mean wind speed field (figure 3a), particularly in the shear layers in the building wakes, where the large shear produces large amounts of TKE. The building cluster area, e.g. at the top right, can be seen to have low levels of turbulence.
@@
Figure 4: Plane-averaged flow statistics. (a) streamwise velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ against normalised height $z/h_{m}$, overlaid with logarithmic profiles (33) estimated using the log-law fitting and the Jackson displacement method; (b) normalised kinematic turbulent shear stress $-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and dispersive stress $-\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$. The dotted line marks the maximum building height $h_{\max}$.
@@
Figure 4(a) shows the profile of plane-averaged superficial streamwise velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ with the height and normalised by the friction velocity $u_{\tau}=\sqrt{{\tau_{0}}}$. Within the canopy region, i.e., $z\leq h_{\max}$, somewhat surprisingly, the streamwise velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ increases roughly linearly with the height, which is due to the morphology used here. Above the canopy region, the profile fits the logarithmic law
$\left\langle{\smash{\overline{u}}}\right\rangle_{xy}(z)=\frac{u_{\tau}}{\kappa}\ln{\left(\frac{z-z_{d}}{z_{0}}\right)}\,,$ (33)
@@
where $\kappa$ is the Von Karman constant, $z_{d}$ is the displacement length $z_{d}$ and $z_{0}$ is the roughness length. Two different methods will be used to estimate the parameters $z_{d}$ and $z_{0}$. The first is to find the $z_{d}$ and $z_{0}$ directly from the profile in the inertial sublayer region (for example here, we choose the fitted region above $2h_{\max}$ to the top end) that fits the logarithmic law best. This method takes $\kappa=0.41$ to be given and results in $z_{d}=40.0\textnormal{m}\approx 1.33h_{m}$ and $z_{0}=4.5\textnormal{m}\approx 0.15h_{m}$. The second method follows a hypothesis by Jackson (1981), that assumes the displacement height $z_{d}$ corresponds to the centre of pressure of the forces acting on the buildings, i.e.
@@
where $\left\langle{\smash{f_{D}}}\right\rangle_{xy}$ is the distributed drag force and note that $\tau_{D}=\tau_{D;\infty}$ (see §2.5). The displacement length calculated is $z_{d}=35.1\,\textnormal{m}\approx 1.17h_{m}$. In this approach, $\kappa$ is determined together with the roughness length $z_{0}$ to best fit the logarithmic law, with results $\kappa=0.37$, and $z_{0}=6.8\,\textnormal{m}\approx 0.23h_{m}$. The velocity profiles fitted using both methods are plotted in figure 4(a) as the dashed lines. A prediction of $z_{d}$ and $z_{0}$ based on the plane index $\lambda_{p}$ (Macdonald et al. 1998; Kanda et al. 2013) results values $z_{d}\approx 1.20h_{m},z_{0}\approx 0.15h_{m}$ for our case. All our results are generally close to this prediction with $z_{0}$ from the second method slightly higher.
@@
Figure 4(b) shows the vertical evolution of the plane-averaged kinematic turbulent shear stress $-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and dispersive stress $-\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$. It shows that the turbulent shear stress $-\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ increases with height to a peak near the maximum building height, and then decreases linearly with height as expected from classical boundary layer theory. The dispersive flux $-\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$ is significant within the canopy with a peak near the mean building height. Above the canopy, the dispersive stress reduces to a very low value, and we thus conclude that the roughness sublayer for this case extends up to $z=h_{\max}$. We note that the roughness sublayer being identical to the maximum building heights is quite unusual. Indeed, for uniform or random-height cube arrays (Coceal et al. 2006; Xie et al. 2008; Sützl et al. 2020; Lu et al. 2024), the roughness sublayer extends to above the building heights. This is likely the result of having a very large domain with a specific non-repeating height distribution of buildings. There are 39 buildings out of the total 512 that have height $h_{\max}$, which is apparently insufficient to induce significant perturbations to the mean velocity field.
@@
Figure 5: Plane-averaged momentum and cumulative stresses as a function of normalised height $z/h_{m}$. (a) Momenum balance (21). (b) Cumulative kinematic stress balance (25). Dotted lines indicate the maximum building height.
@@
Figure 5(a) shows the plane-averaged force balance. The drag force $\left\langle{\smash{f_{D}}}\right\rangle_{xy}$ is determined in two ways: from the residual of the other two terms and from its direct definition (16). The good agreement between both shows that the momentum budget closes and is thus confirmation that the system is in a statistically steady state. Within the canopy region, the magnitude of drag force decreases with height as expected because less space is occupied by building areas at larger heights. Above the canopy, the drag force vanishes and the pressure gradient matches the turbulent shear stress divergence. Figure 5(b) illustrates the plane-averaged kinematic stress balance — the integral version of the force balance (25). Above the canopy region, the stresses $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ and $\tau_{f}$ coincide in the absence of the drag force and both are linear with the height due to the constant gradient shown in figure 5(a). Within the canopy region, the kinematic shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ increases roughly linearly with height, and the cumulative drag stress $\tau_{D}$ reduces from its maximal value at $z=0$ to zero at $h_{\max}$.
@@
There are two methods in which the drag force is typically parameterised. The first is to introduce a height-dependent drag coefficient $C_{d}(z)$ related to the forcing as (Santiago and Martilli 2010)
@@
$\zeta(z)=A_{F}^{-1}\int_{z}^{h_{\max}}L_{b}(z^{\prime})\textnormal{d}z^{\prime}\,.$ (37)
@@
Figure 6(b) shows the drag coefficient $C_{d}$ calculated from (35). The coefficient is the largest near the ground due to very low velocities. The coefficient decreases from $4.0$ to $0.5$ over most of the building heights, i.e., $z/h_{\max}\geq 0.1$. The coefficient profile is generally consistent with the values reported in Coceal and Belcher (2004). Figure 6(c) shows the relation between $\tau_{D}$ and $\zeta$ together with (36). There is reasonably good agreement of the data with the parameterisation, although the current simulation does not show an inflection point near ground level ($\zeta=0)$. This might suggest that (36), which was obtained from a best fit based on eight independent simulations may need some further calibration, but this would need many simulations (e.g. using the Lu et al. (2024) dataset) to be done reliably. We will therefore keep using (36) even though we could construct a better fitting relation.
Figure 6: Distributed drag parameterisations as a function of scaled height $z/h_{\max}$. (a) Total frontal building width $L_{b}$ and scaled total frontal building area $\zeta$; (b) drag coefficient $C_{d}$ calculated from (35); (c) plane-averaged normalised cumulative drag stress $\tau_{D}/\tau_{0}$ against $\zeta$, overlaid with parameterisation (36). The dotted line indicates the mean building height.
@@ 4.2 Multi-resolution analysis
In this section, we will use the multi-resolution framework developed in §2. To illustrate the coarse-graining, the mean streamwise velocity $\overline{u}$ at the mean building height $z=30$ m is shown in figure 7 at a number of different averaging lengths $L$. Figure 7(a) shows the velocity field at the original resolution. Recall that the resolution is $1.5$ m in both directions and that the white boxes represent the buildings where there is no fluid. Figure 7(b-e) show the superficial velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{L}$ at averaging lengths $L$ of 6, 24, 96 and 384 m. Figures 7(b,c) show that the superficial averaging gradually fills in the areas in which no information about the flow is present. Figure 7(c) shows that the values of $\left\langle{\smash{\overline{u}}}\right\rangle_{24}$ are unrealistically low inside the buildings, as can be expected theoretically (Whitaker 1999; Schmid et al. 2019). At $L=96$ m however (figure 7d), this no longer appears to be a significant effect. At a resolution $L=384$ m (figure 7e), the velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{384}$ is practically homogeneous. Figure 7(f) shows the plane averaged value $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ of the plane, which is homogeneous by definition.
Figure 8: The total, resolved and unresolved variances of the streamwise velocity $\overline{u}$ (top panels) and the TKE $k$ (bottom panels) as a function of the averaging length $L$. The variances are taken at three different heights within the canopy region: (a,d) $z/h_{m}=0.5$, (b,e) the mean building height $z/h_{m}=1.0$, (c,f) the canopy top $z/h_{m}=2.5$.
The variance of any quantity $\overline{\varphi}$ can be used to quantify its spatial heterogeneity (see also Yu et al. 2023). The plane-averaged variance of $\overline{\varphi}$ is defined as
$\sigma^{2}(z)=\left\langle{\smash{(\overline{\varphi}(\boldsymbol{x}_{\perp},z)-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})^{2}}}\right\rangle_{xy}\,,$ (38)
@@
where the unresolved variance $\sigma^{2}_{U}$, resolved variance $\sigma^{2}_{R}$ and (twice) the interaction covariance ${\sigma^{2}_{I}}$ are defined as, respectively,
@@
${\sigma^{2}_{U}}$ $={\left\langle{\smash{(\overline{\varphi}-\left\langle{\smash{\overline{\varphi}}}\right\rangle)^{2}}}\right\rangle_{xy}},$
@@
${\sigma^{2}_{R}}$ $={\left\langle{\smash{(\left\langle{\smash{\overline{\varphi}}}\right\rangle-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})^{2}}}\right\rangle_{xy}},$
@@
${\sigma^{2}_{I}}$ $={2\left\langle{\smash{(\overline{\varphi}-\left\langle{\smash{\overline{\varphi}}}\right\rangle)(\left\langle{\smash{\overline{\varphi}}}\right\rangle-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})}}\right\rangle_{xy}}.$
These variances are dependent on the height $z$ and averaging length $L$.
@@
Figure 8 shows the plane-averaged variances of streamwise velocity $\overline{u}$ and TKE $k$ as a function of the averaging length $L$ at three levels within the canopy region, specifically, near the ground, at the mean height and at the maximum building height. The velocity and TKE variance data have very similar trends. By definition, the total variance $\sigma^{2}$ is independent of the averaging length $L$. Generally for both quantities, the variances are small near the ground and are larger at the mean height and the maximum height levels due to the building height distribution. As is visible in figures 3(c) and (f), above the canopy region the total variance vanishes as the flow becomes uniform.
@@
The dependence of the variances on the averaging length $L$ is similar for all heights and both quantities. The resolved variances $\sigma^{2}_{R}$ decrease while the unresolved variances $\sigma^{2}_{U}$ increase with increasing $L$. This is consistent with figure 7, which shows inhomogeneity (large resolved variance) at low values of $L$ and homogeneity (low resolved variance) at high values of $L$, explaining the resolved variance.
@@
The interaction term $\sigma^{2}_{I}$ is always negligible at very high or low values of $L$ as either $\sigma^{2}_{R}$ or $\sigma^{2}_{U}$ would dominate in (39). Indeed, for very high resolutions, most variance is resolved, i.e., it is not necessary to model the unresolved part; while for very low resolutions, the variance is largely unresolved. The interaction term thus provides information on the scales at which resolved and unresolved scales interact most strongly, and thus provides information on the characteristic urban length scale $\ell$ introduced in §2. For all figures, $\sigma^{2}_{I}$ peaks around $L\approx 24-96$ m, and therefore $\ell\approx 50$ m for this case, which is about twice the building width $W$. Note that this value is consistent with the averaging length at which the resolved variance $\sigma^{2}_{R}$ and unresolved variance $\sigma^{2}_{U}$ are equal to each other, which would be an alternative method to infer a characteristic urban lengthscale $\ell$. For averaging lengths $L>200$ m, most variance is unresolved and homogeneity can be assumed.
@@
Figure 9: Vertical profile of plane-averaged cumulative drag stress $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ in the canopy layer against $\zeta$. (a) $L=96$ m. (b) $L=192$ m. (c) $L=384$ m. (d) $L=786$ m, The standard deviation $\sigma_{R}$ is marked by the red band.

### v2

@@ 4 Results
4.1 Classical plane-averaged analysis
@@
Figure 3: Time-averaged flow statistics. (a) Wind speed $\sqrt{\overline{u}^{2}+\overline{v}^{2}}$ at mean building height $z=30$ m overlaid with $(\overline{u},\overline{v})$ velocity vectors; (b) wind speed $\sqrt{\overline{u}^{2}+\overline{w}^{2}}$ at the central vertical plane $y=384$ m overlaid with ($\overline{u}$$\overline{w}$) velocity vectors; (c) Turbulent kinetic energy $k$ at mean building height $z=30$ m. (d) Turbulent kinetic energy $k$ at the central vertical plane $y=384$ m. The white boxes represent the buildings.
@@
The time-averaged wind speed at the mean building height $z=30$ m (Fig. 3a) shows large variations in wind speed and channelling in regions where buildings are absent. A strong reduction in wind speed is shown near some of the tall building clusters (Mishra et al. 2023, 2024), with wakes clearly visible downstream of the tall buildings. The elevation view (Fig. 3b) shows the general increase of wind speed with height and also demonstrates the homogenisation of the wind in the $x$ direction with increasing height, consistent with the extent of the roughness sublayer (the height of this layer is $h_{\max}$ as will be shown in Fig. 4). The velocity vectors show a clearly visible wake behind the building (in particular behind the first building in Fig. 3b). Accelerations are visible near the upwind top of the building (e.g., for the second cube in Fig. 3b), which is consistent with Coceal et al. (2007). In Fig. 3(b), the skimming flow can be observed in places, e.g., near the last cube; and also the flow recirculation, e.g., between the last third and last second cube.
@@
Figures 3(c, d) show the plane view at $z=30$ m and the vertical view at the central plane of time-averaged turbulent kinetic energy (TKE, $k(\boldsymbol{x})\equiv\frac{1}{2}\left(\overline{u^{\prime}u^{\prime}}+\overline{v^{\prime}v^{\prime}}+\overline{w^{\prime}w^{\prime}}\right)$), respectively. Note that the TKE field is correlated with the mean wind speed field: in the building cluster area, e.g. at the top right of Fig. 3(c), it can be seen to have low levels of turbulence due to the low velocity in the building wake; while at the upwind top of the building (e.g., the second building in Fig. 3d), large TKE can be observed due to the acceleration and large shear at the top.
@@
Figure 4: Plane-averaged flow statistics. (a) streamwise velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ against normalised height $z/h_{m}$, overlaid with logarithmic profiles Eq. (33) estimated using the log-law fitting and the Jackson displacement method; (b) normalised kinematic turbulent shear stress $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and dispersive stress $\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$. The dotted line marks the maximum building height $h_{\max}$.
@@
Figure 4 presents the streamwise velocity and turbulent transport, as the pressure gradient forcing is applied solely in the streamwise direction, which governs the dominant flow and transport. Figure 4(a) shows the profile of plane-averaged superficial streamwise velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ and normalised by the friction velocity $u_{\tau}=\sqrt{{\tau_{0}}}$. Above the canopy region i.e., $z\geq h_{\max}$, the profile fits the logarithmic law
[…]
@@
where $\kappa$ is the Von Karman constant, $z_{d}$ is the displacement length and $z_{0}$ is the roughness length. Two different methods will be used to estimate the parameters $z_{d}$ and $z_{0}$. The first is to find the $z_{d}$ and $z_{0}$ directly from the profile in the inertial sublayer region (for example, here, we choose the fitted region above $2h_{\max}$ to the top end) that fits the logarithmic law best. This method takes $\kappa=0.41$ to be given and results in $z_{d}=40.0\textnormal{m}\approx 1.33h_{m}$ and $z_{0}=4.5\textnormal{m}\approx 0.15h_{m}$. The second method follows a hypothesis by Jackson (1981), that assumes the displacement height $z_{d}$ corresponds to the centre of pressure of the forces acting on the buildings, i.e.
@@
where $\left\langle{\smash{f_{D}}}\right\rangle_{xy}$ is the distributed drag force and note that $\tau_{D}=\tau_{D;\infty}$ (see §2.5). The displacement length calculated is $z_{d}=35.1\,\textnormal{m}\approx 1.17h_{m}$. In this approach, $\kappa$ is determined together with the roughness length $z_{0}$ to best fit the logarithmic law, with results $\kappa=0.37$, and $z_{0}=6.8\,\textnormal{m}\approx 0.23h_{m}$. The velocity profiles fitted using both methods are plotted in Fig. 4(a) as the dashed lines. A prediction of $z_{d}$ and $z_{0}$ based on the plane index $\lambda_{p}$ (Macdonald et al. 1998; Kanda et al. 2013) results values $z_{d}\approx 1.20h_{m},z_{0}\approx 0.15h_{m}$ for our case. All our results are generally close to this prediction with $z_{0}$ from the second method slightly higher. As the displacement height $z_{d}$ and roughness length $z_{0}$ is typically related to the morphological parameters like plane area index $\lambda_{p}$, frontal area index $\lambda_{f}$, mean building height $h_{m}$ and the variance of the building height (Grimmond and Oke 1999) which are all predetermined for this case according to the London morphology profile, so that in general, we do not except a big variation in $z_{d}$ and $z_{0}$ among the random building-height distributions, although slight discrepancy would exist for example a distribution with a denser high-rise area may reduce the roughness length due to the sheltering effect (Oke et al. 2017).
@@
Figure 4(b) shows the vertical evolution of the plane-averaged kinematic turbulent shear stress $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ and dispersive stress $\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$. It shows that the turbulent shear stress $\left\langle{\smash{\overline{w^{\prime}u^{\prime}}}}\right\rangle_{xy}$ increases with height to a peak near the maximum building height, and then decreases linearly with height as expected from classical boundary layer theory. The dispersive flux $\left\langle{\smash{\overline{w}^{\prime\prime}\overline{u}^{\prime\prime}}}\right\rangle_{xy}$ is significant within the canopy with a peak near the mean building height. Above the canopy, the dispersive stress reduces to a very low value, which indicates that the upper limit of the roughness sublayer, or equivalently, the lower limit of the inertial sublayer, is at about $z=h_{\max}$. Note that this is a simple approximation; use of other quantities or definitions for the roughness sublayer, such as that of Placidi and Ganapathisubramani (2015), might work better for varying building heights but are beyond the scope of this study. The fact that the upper limit of the roughness sublayer is identical to the maximum building heights is quite unusual, since it implies that there is only an urban canopy layer but no roughness sublayer. Indeed, for uniform or random-height cube arrays (Coceal et al. 2006; Xie et al. 2008; Sützl et al. 2020; Lu et al. 2024), the roughness sublayer extends above the building heights. This is likely the result of having a very large domain with a specific non-repeating height distribution of buildings. There are 39 buildings out of the total 512 that have height $h_{\max}$, which is apparently insufficient to induce significant perturbations to the mean velocity field.
@@
Figure 5: Plane-averaged momentum and cumulative stresses as a function of normalised height $z/h_{m}$. (a) Momenum balance Eq. (21). (b) Cumulative kinematic stress balance Eq. (25). Dotted lines indicate the maximum building height.
@@
Figure 5(a) shows the plane-averaged force balance. The drag force $\left\langle{\smash{f_{D}}}\right\rangle_{xy}$ is determined in two ways: from the residual of the other two terms and from its direct definition Eq. (16). The good agreement between both shows that the momentum budget closes and is thus confirmation that the system is in a statistically steady state. Within the canopy region, the magnitude of the drag force decreases with height as expected because less space is occupied by building areas at larger heights. Above the canopy, the drag force vanishes and the pressure gradient matches the turbulent shear stress divergence. Figure 5(b) illustrates the plane-averaged kinematic stress balance — the integral version of the force balance Eq. (25). Above the canopy region, the stresses $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ and $\tau_{f}$ coincide in the absence of the drag force and both are linear with the height due to the constant gradient shown in Fig. 5(a). Within the canopy region, the kinematic shear stress $\left\langle{\smash{\tau_{z}}}\right\rangle_{xy}$ increases roughly linearly with height, and the cumulative drag stress $\tau_{D}$ reduces from its maximal value at $z=0$ to zero at $h_{\max}$.
@@
We consider two methods in which the drag force can be parameterised. The first is to introduce a height-dependent drag coefficient $C_{d}(z)$ related to the forcing as (Santiago and Martilli 2010)
@@
$\zeta(z)\equiv A_{F}^{-1}\int_{z}^{h_{\max}}L_{b}(z^{\prime})\textnormal{d}z^{\prime}\,.$ (37)
@@
Figure 6(b) shows the drag coefficient $C_{d}$ calculated from Eq. (35). The coefficient is the largest near the ground due to very low velocities. The coefficient decreases from $4.0$ to $0.5$ over most of the building heights, i.e., $z/h_{\max}\geq 0.2$. The coefficient profile is generally consistent with the values reported in Coceal and Belcher (2004). The kink in the $C_{d}$ near $0.1<z/h_{\max}<0.15$ is due to the rapid decreases in building width $L_{b}$, because due to Eq. (35) and drag and velocity profile, the production of $C_{d}$ and $L_{b}$ should decrease smoothly with the height. Figure 6(c) shows the relation between $\tau_{D}$ and $\zeta$ together with Eq. (36). There is reasonably good agreement of the data with the parameterisation, although the current simulation does not show an inflection point near ground level ($\zeta=1)$. This might suggest that Eq. (36), which was obtained from a best fit based on eight independent simulations, may need some further calibration, but this would need many simulations (e.g. using the Lu et al. (2024) dataset) to be done reliably. We will therefore keep using Eq. (36) even though we could construct a better fitting relation.
Figure 6: Distributed drag parameterisations as a function of scaled height $z/h_{\max}$. (a) Total frontal building width $L_{b}$ and scaled total frontal building area $\zeta$; (b) drag coefficient $C_{d}$ calculated from Eq. (35); (c) plane-averaged normalised cumulative drag stress $\tau_{D}/\tau_{0}$ against $\zeta$, overlaid with parameterisation Eq. (36). The dotted line indicates the location of the mean building height.
@@ 4.2 Multi-resolution analysis
In this section, we will use the multi-resolution framework developed in §2. To illustrate the coarse-graining, the mean streamwise velocity $\overline{u}$ at the mean building height $z=30$ m is shown in Fig. 7 at a number of different averaging lengths $L$. Figure 7(a) shows the velocity field at the original resolution. Recall that the resolution is $1.5$ m in both directions and that the white boxes represent the buildings where there is no fluid. Figure 7(b-e) show the superficial velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{L}$ at averaging lengths $L$ of 6, 24, 96 and 384 m. Figures 7(b,c) show that the superficial averaging gradually fills in the areas in which no information about the flow is present. Figure 7(c) shows that the values of $\left\langle{\smash{\overline{u}}}\right\rangle_{24}$ are unrealistically low at the original building locations, as can be expected theoretically (Whitaker 1999; Schmid et al. 2019). At $L=96$ m however (Fig. 7d), this no longer appears to be a significant effect. At a resolution $L=384$ m (Fig. 7e), the velocity $\left\langle{\smash{\overline{u}}}\right\rangle_{384}$ is practically homogeneous. Figure 7(f) shows the plane-averaged value $\left\langle{\smash{\overline{u}}}\right\rangle_{xy}$ of the plane, which is homogeneous by definition.
Figure 8: The total, resolved and unresolved variances of the streamwise velocity $\overline{u}$ (top panels) and the TKE $k$ (bottom panels) as a function of the averaging length $L$. The variances are taken at three different heights within the canopy region: (a,d) the pedestrian level $z/h_{m}=0.05$, (b,e) the mean building height $z/h_{m}=1.0$, (c,f) the canopy top $z/h_{m}=2.5$.
The variance of any quantity $\overline{\varphi}$ can be used to quantify its spatial heterogeneity (see also Yu et al. 2023). The plane-averaged variance of $\overline{\varphi}$ quantifies the deviation of the original field from its plane average, which is defined as
$\sigma^{2}(z)\equiv\left\langle{\smash{(\overline{\varphi}(\boldsymbol{x}_{h},z)-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})^{2}}}\right\rangle_{xy}\,$, (38)
@@
where the unresolved variance $\sigma^{2}_{U}$ quantifies the deviation of the original field from the spatial average, the resolved variance $\sigma^{2}_{R}$ quantifies the deviation of the spatial average from the plane average, and (twice) the interaction covariance $\sigma^{2}_{I}$} are defined as, respectively,
@@
${\sigma^{2}_{U}}$ $\equiv{\left\langle{\smash{(\overline{\varphi}-\left\langle{\smash{\overline{\varphi}}}\right\rangle)^{2}}}\right\rangle_{xy}},$
@@
${\sigma^{2}_{R}}$ $\equiv{\left\langle{\smash{(\left\langle{\smash{\overline{\varphi}}}\right\rangle-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})^{2}}}\right\rangle_{xy}},$
@@
${\sigma^{2}_{I}}$ $\equiv{2\left\langle{\smash{(\overline{\varphi}-\left\langle{\smash{\overline{\varphi}}}\right\rangle)(\left\langle{\smash{\overline{\varphi}}}\right\rangle-\left\langle{\smash{\overline{\varphi}}}\right\rangle_{xy})}}\right\rangle_{xy}}.$
These variances are dependent on the height $z$ and the averaging length $L$.
@@
Figure 8 shows the plane-averaged variances of streamwise velocity $\overline{u}$ and TKE $k$ as a function of the averaging length $L$ at three levels within the canopy region, specifically, at the pedestrian level, at the mean height and at the maximum building height. The velocity and TKE variance data have very similar trends. By definition, the total variance $\sigma^{2}$ is independent of the averaging length $L$. Generally, for both quantities, the variances are small near the ground, are larger at the mean height and the maximum height levels due to the building height distribution, and should gradually vanish above the canopy region as the flow becomes uniform.
@@
The dependence of the variances on the averaging length $L$ is similar for all heights and both quantities. The resolved variances $\sigma^{2}_{R}$ decrease while the unresolved variances $\sigma^{2}_{U}$ increase with increasing $L$. This is consistent with Fig. 7, which shows inhomogeneity (large resolved variance) at low values of $L$ and homogeneity (low resolved variance) at high values of $L$, explaining the resolved variance.
@@
The interaction term $\sigma^{2}_{I}$ is always negligible at very high or low values of $L$ as either $\sigma^{2}_{R}$ or $\sigma^{2}_{U}$ would dominate in Eq. (39). Indeed, for very high resolutions, most variance is resolved, i.e., it is not necessary to model the unresolved part; while for very low resolutions, the variance is largely unresolved. The interaction term thus provides information on the scales at which resolved and unresolved scales interact most strongly, and thus provides information on the characteristic urban length scale $\ell$ introduced in §2. For all figures, $\sigma^{2}_{I}$ peaks around $L\approx 24-96$ m, and therefore $\ell\approx 50$ m for this case, which is about twice the building width $W$. Note that this value is consistent with the averaging length at which the resolved variance $\sigma^{2}_{R}$ and unresolved variance $\sigma^{2}_{U}$ are equal to each other, which would be an alternative method to infer a characteristic urban lengthscale $\ell$. For averaging lengths $L>200$ m, most variance is unresolved and homogeneity can be assumed.
@@
Figure 9: Vertical profile of plane-averaged cumulative drag stress $\left\langle{\smash{\tau_{D;L}}}\right\rangle_{xy}$ in the canopy layer against the normalised frontal area $\zeta$. (a) $L=96$ m. (b) $L=192$ m. (c) $L=384$ m. (d) $L=786$ m, The standard deviation $\sigma_{R}$ is marked by the red band.